OpenStudy (anonymous):

Evaluate 3 to the power of -1

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) the negative power puts the original number under a 1 so for example 5^(-1) = 1/(5^1) = 1/5 4^(-2) = 1/(4^2) = 1/16 etc do you understand this? :3

OpenStudy (abdullah1995):